UT 20140117 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: SAllenson
LBTO Support: TSargent
Visitor: CVeillet, MWagner, JBrynnel, DSummers

Plan for the night

- put in AGw-transform and check it
- Photometric gradient (need to check weather before starting)
- Longslit-Acq-tests
- MOS-acq tests
- Distortion determination, image scale
- Image movement due to flexure

Summary

GEIRS ran the entire night without any failure!

SW was also very stable, except one repeated failure with the telescope service

AGw-transform finally works

MOS-acquisition works great

LS_acquisition with 2 stars works in principle (stars are in slit), but has offset in y by 10-40pix wrt to the requested position

Took further throughput, center of rotation, distortion, photometric gradient data

Took W3 data in H+J

Preparations

- A visiting team come up to see how it goes (Christian, Mark, Joar & Doug)
- Right after flats a new AGw-transform based on last nights data will be put in
- Checked field alignment (off by dx=1200 steps, dy = -300steps)

Details

All times are UT, files are luci2.20140116.iiii

Twilight flats

00:28, Br_gam, telescope in zenith
#15-24, Br_gam, 10*3sec, 10k
#25-34, Br_gam, 10*6sec, 20k
#35-44, Br_gam, 10*2.7sec, 8k
#45-54, Br_gam, 10*2.7sec, 5k
00:45
#55-64, OH_1060, 10*2.7sec, 28k
#65-74 (#74 crap), P_gam, 10*2.7sec, 24k
#75-76, Br_gam, 10*2.7sec, 2k, (done erroneously)
#77-86, OH_1060, 10*2.7sec, 10k
#87-96, P_gam, 10*2.7sec, 7k
#97-106, OH_1060, 10*2.7sec, 2k
#107-116, P_gam, 10*2.7sec, 1k
01:05
#117-126, J, 10*7sec, 20k -> 11k
#127-127, J, 10*7sec, 10k -> 8k

00:40 Sunset, weather: clear, some cirrus down in NW which may come in later on
may probably not an issue for flats, low wind, T=1deg, hum=20%

AGw-check, LS-acquisition

01:10 Steve checks pointing & collimation on BS9107

01:21 Point to BS9107 for LSacq & AGw-check, PA=0, 1.5" slit, sky #137, src #138, LS #139
verification #140, OK, offset -20" #141, look good, offset -20", #142, looks good
offset +60", #143, offset +20", looks still good

01:51 BS9107, PA=225, recheck AGw, #144-150, OK

AGw-transform works

02:07 Point to BS9107 for LS-acq 2 stars, PA=0, 1.5" slit, sky, src, slit, #151-153
verification #154, alignment OK, but stars are about 20pix off in y wrt to the requested position

02:22 Point to BS9107 for LS-acq 2 stars, PA=225, 1.5" slit, sky, src, slit, #155-157
verification #158, alignment OK, offset 50pix in y,

02:39 Photometric gradient, in Br_gam & Ks, redo from November, #158-178 , seing 0.5-0.6"
stopped at the last Br_gam exp., restarted LUCI2 execution panel, OK, reason unclear
, probably due to Alexander who tried to work on other issues at the same time

NXwindows crashed (visiting team to blame!!!), restartet NXClient OK

W3, nice images for LUCI-promotion

03:26 W3, seeing is good, will try to collect good JH-data, 1st obsitem, acq. in K-band, #179, 0.55" seeing
H #180-213 (stopped after exp 211, commit not submitted to telescope, same problem as last night
stop/restart telescope service and continue), seeing typically 0.6-0."8
J #214-247, seeing degraded considerably during J-band exp. from 0.6-0.7" - 1",
probably not very useful... Had to stop/restart telescope service after #243
Images #248-253 crap

Distortion determination

05:42 M67...too close to the moon, not tonight...

Quite some cirrus coming in

Center of rotation

05:50 BS9126, PA=0, 90, 180, 270, #254-261

Center of rotation: x,y = 976,1063

Distortion determination

06:22 Michelle modified her script, N3.75 #262-270, N1.8 #271-279, seeing 0.7"-1.5"
Had to stop/restart telescope service after ima #272

Cirrus have mostly passed

MOS-acquisition

06:55 Stone-field sl01, needed to modify script by scripting tool, had to replace in tag maskID
the ID-number (ID943523) by mask-name (sl01)
Done with N1.8 camera: sky #280, src #281, MOS-view #282, alignment OK, need to modify pos of some of the stars since its crowded field
and sometimes other brighter ones were in the searching region
confirmation #283, looks very good

07:42 Stone-field sl03, script also modified, done with N3.75-camera
sky #284, src #285, MOS-view #286/287, had to stop/start telescope service after #284
alignment easy, all stars found and well centered, reason is search radius in pixel
confirmation #288, SUPER!!!

08:00 Mirror panicked, some work for Steve, collimated on a nearby star

08:13 Stone-field sl01, with N3.75, sky #289, src #290, MOS-view #291
alignment again easy, verification #292, SUPER!!!
Time from start of the script (moving target to source) until alignment is verified 12min!!!

08:30 M67, close to moon, lets give it a try, too close

Throughput - since offsets work now

09:12 BS9142 in zJHKKS for throughput-calib, #299-306 stop/restart telescope service after #300

AI Alexander/Tom get rid of problem with telescope service

09:30 Throughput BS9142_all_corrected, spectra #313- #340

10:00 Several problems with Telescope Service during throughput measurements, two solved with stop/restart but the most severe, after image 320, required more work. Had to restart several times and then restart the Observation Manager

10:54 stop/restart telescope service after 334; noticed filter wheel one was also "stopped" in the software manager and turned it "on"

Same setup as in previous nights.

11:00 Start telluric for BS9142, HIP3735. Some problems with the number of exposures. Fixed scripts and continued. Relevant acq files are 351, 352, 353

11:32 Science started. Telluric 355-368. Stop/start telescope service again before 358

LS/MOS-tests

11:50 BS9148, check for alignment of 2 stars in slit, #369 src, #370 slitview, #371 verification, #372 verification wo slit, works in principle, but stars are offset in y
wrt to requested position.
Pos stars: 965.3/1069.9 // 902.6/777
Pos slit: 1013.8/1023 // 1018.1/722.9
Required: dx,dy=6.9",-3.8",-12.9
Stars found: 1024/1038 // 1028/739
Stars are in slit, BUT offset by 15pix in y wrt to the requested position

12:15 SDSS1110, JKurks mask with curved slits, #373 sky, #374 src, #375 MOS
aligned, #376 verification, PERFECT

MOS alignment really works super & fast

Tests big offsets - NGC 5053

12:43 NGC5053_dither, J and KS, problems with collimation, Steve restartet GCS.
Do NGC5053 in Ks only, #388-398 , J-band images taken before are probably not useful

13:52 Stopped, too bright for guider, Steve brings telescope cose to zenith & tracking

Twilight flats

Try to do it in track mode and offset 10" in between each flat via tel.service.gui, tel @ 86 ele

J, #399-405, 3sec, 13k - 33k
14:00
#406 crap
K, #407-424, 3sec, 9k - 30k
14:11
Br_gam #425-434 3sec, , 4-5k
#435-444, 6sec, 10-15k
#4445-454, 8sec, 21-27k

14:25 Closing

Daytime calibrations

UT 20140117
Spectroscopic calibrations
/home/lucifer/COMMISSIONING_STUFF_ONLY/Jan2014_Calib
manually moving FW2 to next filter to avoid problems

G210H_SpecFlats.xml
Ran OK with Alexander's change
Problem stabilizing grating @1.65um
  1. k ADU with Halo1

G210z_SpecFlats.xml
My change first time was incorrect
Corrected, commit 2 worked
  1. K ADU in 30s with Halo1
    Dichroic cutoff about in middle of detector!
    Should use different CWL for z spectra?

G210J_SpecFlats.xml
Problem stabilizing at 1.25um
Manual re-request of 1.25um worked
  1. k ADU in 5.0s

G210K_SpecFlats.xml
  1. K ADU in 2.75s

Note: Halo1 about 3x as bright as in the acceptance test report.

G200zJ_SpecFlats.xml
  1. 75s w/Halo1 34K ADU
  2. s w/Halo2 14K ADU
    Edge of dichroic cut-on visible @ left edge...change CWL?
    Countdown timer on READ2 gui going to -3 by end of exposure
    Halo2 may have also been on for these.

G200HK_SpecFlats.xml
  1. 75s w/Halo1 saturated
  2. 0s w/Halo2 10k ADU 2k ADU/sec

Note: Halo2 about 2x as bright as in the acceptance test report.

Check if both FC dots on ACTFLEX_3LS are visible...
Lower one there but much fainter than one at top...hole blocked?

G150KsN180_SpecFlats.xml
N1.8 Ks 2.17um Halo2 2.75s=3.3k 10.0s=11k ADU
Spectrum offset to left of center...need different CWL to center on N375?

G150KsN375_SpecFlats.xml
Halo1 2.75s 18k ADU
Spectrum offset to left of center...need different CWL to center on N375?

Check if files are saved when taken from OBs when "Manually save last
exposure..." is checked on READ2 GUI.

-- JochenHeidt - 17 Jan 2014
Topic revision: r10 - 19 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback